# Mitigating Suboptimality of Deterministic Policy Gradients in Complex Q-functions

### v1

@@ C.2 RecSim
The click probability is computed with a softmax over the item score and a predefined skip score:
$p_{item}$ $=\frac{e^{\text{score}_{item}}}{e^{\text{score}_{item}}+e^{\text{score}_{skip}}},\quad p_{skip}=1-p_{item}$
The user then stochastically chooses to click or skip based on this distribution.
Action Representations: Following Jain et al. (2021), items are represented as continuous vectors sampled from a Gaussian Mixture Model (GMM), with centers representing item categories.
C.3 Continuous Control
MuJoCo (Todorov et al., 2012) is a physics engine that provides a suite of standard reinforcement learning tasks with continuous action spaces, commonly used for benchmarking continuous control algorithms. We briefly describe some of these tasks below:
Hopper: The agent controls a one-legged robot that must learn to hop forward while maintaining balance. The objective is to maximize forward velocity without falling.
Walker2d: The agent controls a two-legged bipedal robot that must learn to walk forward efficiently while maintaining balance. The goal is to achieve stable locomotion at high speeds.
HalfCheetah: The agent controls a planar, cheetah-like robot with multiple joints in a 2D environment. The task requires learning a coordinated gait to propel the robot forward as quickly as possible.
Ant: The agent controls a four-legged, ant-like robot with multiple degrees of freedom. The challenge is to learn to walk and navigate efficiently while maximizing forward progress.
C.3.1 Restricted Locomotion in Mujoco
The restricted locomotion Mujoco tasks are introduced to demonstrate how common DPG-based approaches get stuck in local optima when the Q-landscape is complex and non-convex. This setting limits the range of actions the agent can perform in each dimension, simulating realistic scenarios such as wear and tear of hardware. For example, action space may be affected as visualized in Figure 6. A mixture-of-hypersphere action space is used to simulate such asymmetric restrictions, which affect the range of torques for the Hopper and Walker joints, as well as the forces applied to the inverted pendulum and double pendulum. The hyperspheres are sampled randomly, and their size and radius are carefully tuned to ensure that the action space has enough valid actions to solve the task.
Definition of restriction.
@@ C.3.1 Restricted Locomotion in Mujoco
    Invalid action vectors are replaced with 0 by changing the environment’s step function code:
@@
    Invalid action vectors are replaced with -1 by changing the environment’s step function code:
@@
    The action space is $1$-dimensional, with force applied on the cart. The implication is that the cart is pushed in the left direction for $0.02$ (default) seconds. Note that the action vectors are not zeroed because a 00-action is often the optimal action, particularly when the agent starts upright. This would make the optimal policy trivially be learning to select invalid actions.
Appendix D Additional Results
D.1 Experiments on Continuous Control (Unrestricted Mujoco)
In standard MuJoCo tasks, the Q-landscape is likely easier to optimize compared to MuJoCo-Restricted tasks. In Figure 15, we find that baseline models consistently perform well in all standard tasks, unlike in MuJoCo-Restricted tasks. Thus, we can infer the following:
@@ D.1 Experiments on Continuous Control (Unrestricted Mujoco)
    Baselines have sufficient capacity, are well-tuned, and can navigate simple Q-landscapes optimally.
@@
    SAVO performs on par with other methods in MuJoCo tasks where the Q-landscape is easier to optimize, showing that SAVO is a robust, widely applicable actor architecture.
@@
    Baselines performing well in unrestricted locomotion but suboptimally in restricted locomotion delineates the cause of suboptimality to be the complexity of the underlying Q-landscapes, such as those shown in Figure 2. SAVO is closer to optimal in both settings because it can navigate both simple and complex Q-functions better than alternate actor architectures.
Figure 15: Unrestricted LocomotionD.1). SAVO and most baselines perform optimally in standard MuJoCo continuous control tasks, where the Q-landscape is easy to navigate.
Figure 16: Ablations of SAVO variationsD.2) shows the importance of (i) the approximation of surrogates, (ii) removing TD3’s action smoothing, (iii) conditioning on preceding actions in the successive actor and surrogate networks, and (iv) individual actors that separate the action candidate prediction instead of a joint high-dimensional learning task.
D.2 Per-Environment Ablation Results
@@
Removing information about preceding actions does not significantly degrade SAVO’s performance since preceding actions’ Q-values are indirectly incorporated into the surrogates’ training objective (see Eq. (9)), except for MineWorld where this information helps improve efficiency.
@@
Figure 17: SAC is orthogonal to the effect of SAVOD.3). SAC is a different algorithm than TD3, whereas SAVO is a plug-in actor architecture for TD3. Thus, tasks where SAC outperforms TD3 differ from tasks where SAVO outperforms TD3. Also, TD3 outperforms SAC in Restricted Hopper and Inverted-Double-Pendulum. However, SAVO+TD3 guarantees improvement over TD3.
D.3 SAC is Orthogonal to SAVO
We compare Soft Actor-Critic (SAC), TD3, and TD3 + SAVO across various Mujoco-Restricted tasks. Figure 17 shows that SAC sometimes outperforms and sometimes underperforms TD3. Therefore, SAC’s stochastic policy does not address the challenge of non-convexity in the Q-function. In contrast, SAVO+TD3 consistently outperforms TD3 and SAC, demonstrating the effectiveness of SAVO in complex Q-landscapes. While SAC can be better than TD3 in certain environments, its algorithmic modifications are orthogonal to the architectural improvements due to the SAVO actor.
Figure 18: SAC is suboptimal in complex Q-landscapeD.3) of Restricted Inverted Double Pendulum, but SAVO helps.
In the Restricted Inverted Double Pendulum task (Figure 18), SAC underperforms even TD3. Analogous to TD3, the suboptimality is due to the non-convexity in the soft Q-function landscape, where small changes in nearby actions can lead to significantly different environment returns. We combine SAC with SAVO’s successive actors and surrogates to better maximize the soft Q-function, naively considering action candidates for $\mu_{M}$ as the mean action of the stochastic actors. We observe that this preliminary version of SAC + SAVO shows significant improvements over SAC in complex Q-landscapes. In future work, we aim to formalize a SAVO-like objective that effectively enables SAC’s stochastic actors to navigate the non-convexity of its soft Q-function.
D.4 Increasing Size of Discrete Action Space in RecSim
We test the robustness of our method to more challenging Q-value landscapes in Figure 19, especially in discrete action space tasks with massive action spaces. In RecSim, we increase the number of actual discrete actions from $10,000$ to $100,000$ and $500,000$. The experiments show that SAVO outperforms the best-performing baseline of TD3 + Sampling (Wolpertinger) and the best-performing ablation of SAVO + Joint Action. This shows that SAVO maintains robust performance even as the action space size increases and the Q-function landscape becomes more intricate. In contrast, the baselines experienced performance deterioration as action sizes grew larger.
Figure 19: Increasing RecSim action set sizeD.4). (Left) $100,000$ items, (Right) $500,000$ items (6 seeds) maintains the performance trends of SAVO and the best-performing baseline (TD3 + Sampling) and the best-performing ablation (SAVO with Joint-Action).
Appendix E Validating SAVO Design Choices
E.1 Design Choices: Action summarizers
Three key architectures were considered for the design of the action summarizer: DeepSets, LSTM, and Transformer models, represented by SAVO, SAVO-LSTM, and SAVO-Transformer in Figure 20, respectively. In general, the effect of the action summarizer is not significant, and we choose DeepSet for its simplicity for most experiments.
Figure 20: Action summarizer comparisonE.1). The effect is not significant The results are averaged over 5 random seeds, and the seed variance is shown with shading.
E.2 Conditioning on Previous Actions: FiLM vs. MLP
We examined two approaches for conditioning on the previous action list summary: Feature-wise Linear Modulation (FiLM) and concatenation with input, represented by the FiLM and non-film variants in Figure 21. Across tasks, FiLM outperformed the non-FiLM version, showing the effectiveness of layerwise conditioning in leveraging prior action information for action selection and surrogate value prediction. This shows that the successive actors are appropriately utilizing the actions from the preceding actors to tailor their search for optimal actions, and the successive surrogates can better evaluate Q-values, knowing where they could be thresholded by the loss function.
Figure 21: FiLM to condition on preceding actionsE.2). FiLM ensures layerwise dependence on the preceding actions for acting in actors $\nu_{i}$ and for predicting value in surrogates $\hat{\Psi}_{i}$, which generally results in better performance across tasks.
E.3 Exploration Noise comparison: OUNoise vs Gaussian
(a) Noise types
Figure 22: OU versus Gaussian NoiseE.3). We do not see a significant difference due to this choice, and select OU noise due to better overall performance in experiments.
We compare Ornstein-Uhlenbeck (OU) noise with Gaussian noise across our environments and find that OU noise was generally better, with the difference being minimal. We chose to use OU for our experiments and a comparison on Hopper-Restricted is shown in Figure 22(a). We note that TD3 (Fujimoto et al., 2018) also suggests no significant difference between OU and Gaussian noise and favored Gaussian for simplicity. All our baselines use the same exploration backbone, and we confirm that using OU noise is consistent with the available state-of-the-art results with TD3 + Gaussian noise on common environments like Ant, HalfCheetah, Hopper, and Walker2D.
Appendix F Network Architectures
F.1 Successive Actors
The entire actor is built as a successive architecture (see Figure 4), where each successive actor receives two pieces of information: the current state and the action list generated by preceding actors. Each action is concatenated with the state to contextualize it and then summarized using a list-summarizer, described in §F.3. This list summary is concatenated with the state again and passed into an MLP with ReLU (3 layers for MuJoCo tasks and 4 layers for MineWorld and RecSim) as described in Table 2. This MLP generates one action for each successive actor, which is subsequently used as an input action to the succeeding action lists. For discrete action space tasks, this generated action is processed with a 1-NN to find the nearest exact discrete action. Finally, the actions generated by each individual successive actor are accumulated, and the maximizer actor $\mu_{M}$ step from Eq. (5) selects the highest-valued action according to the Critic Q-network, described in §F.2.
F.2 Successive Surrogates
As Figure 4 illustrates, there is a surrogate network for each actor in the successive actor-architecture. Each successive critic receives three pieces of information: the current state, the action list generated by preceding actors, and the action generated by the actor corresponding to the current surrogate. Each action is concatenated with the state to contextualize it and then summarized using a list-summarizer, described in §F.3. This list summary is concatenated with the state and the current action, and passed into a 2-layer MLP with ReLU (See Table 2). This MLP generates the surrogate value $\hat{\Psi}_{i}(s,a;a_{<i})$ and is used as an objective to ascend over by its corresponding actor $\nu_{i}$.
F.3 List Summarizers
To extract meaningful information from the list of candidate actions, we employed several list summarization methods following Jain et al. (2021). These methods are described below:
Bi-LSTM: The action representations of the preceding actors’ actions are first passed through a two-layer multilayer perceptron (MLP) with ReLU activation functions. The output of this MLP is then processed by a two-layer bidirectional LSTM network (Huang et al., 2015). The resulting output is fed into another two-layer MLP to create an action set summary, which serves as an input for the actor-network (§F.1) and the surrogate network (§F.2).
DeepSet: The action representations of the preceding actors’ actions are initially processed by a two-layer MLP with ReLU activations. The outputs are then aggregated using mean pooling over all candidate actions to compress the information into a fixed-size summary. This summary is passed through another two-layer MLP with ReLU activation to produce the action set summary, which serves as an input for the actor-network (§F.1) and the surrogate network (§F.2).
Transformer: Similar to Bi-LSTM, the action representations of the preceding actors’ actions are first processed by a two-layer MLP with ReLU activations. The outputs are then input into a Transformer network with self-attention and feed-forward layers to summarize the information. The resulting summary is used as part of the input to the actor-network (§F.1) and the surrogate network (§F.2).
F.4 Feature-wise Linear Modulation (FiLM)
Feature-wise Linear Modulation (Perez et al., 2018) is a technique used in neural networks to condition intermediate feature representations based on external information, enhancing the network’s adaptability and performance across various tasks. FiLM modulates the features of a layer by applying learned, feature-wise affine transformations. Specifically, given a set of features $\mathbf{F}$, FiLM applies a scaling and shifting operation,
$\text{FiLM}(\mathbf{F})=\gamma\odot\mathbf{F}+\beta,$
where $\gamma$ and $\beta$ are modulation parameters learned from another source (e.g., a separate network or input), and $\odot$ denotes element-wise multiplication. This approach allows the network to selectively emphasize or de-emphasize aspects of the input data, effectively capturing complex and context-specific relationships. FiLM has been successfully applied in tasks such as visual question answering and image captioning, where conditioning visual features on textual input is essential. We apply FiLM while conditioning the actor and surrogate networks on the summary of preceding actions.
Appendix G Experiment and Evaluation Setup
G.1 Aggregated Results: Performance Profiles
To rigorously validate the aggregate efficacy of our approach, we adopt the robust evaluation methodology proposed by Agarwal et al. (2021). By incorporating their suggested performance profiles, we conduct a comprehensive comparison between our method and baseline approaches, providing a thorough understanding of the statistical uncertainties inherent in our results. Figure 7(a) shows the performance profiles across all tasks. The x-axis represents normalized scores, calculated using min-max scaling based on the initial performance of untrained agents aggregated across random seeds (i.e., Min) and the final performance from Figure 8 (i.e., Max). The results show that our method consistently outperforms the baselines across various random seeds and environments. Our performance curve remains at the top as the x-axis progresses, while the baseline curves decline earlier. This highlights the reliability of SAVO over different environments and 10 seeds.
G.2 Implementation Details

### v2

@@ 11.2 RecSim
$p_{item}$ $=\frac{e^{score_{item}}}{e^{s_{item}}+e^{score_{skip}}}$
$p_{skip}$ $=\frac{e^{score_{skip}}}{e^{s_{item}}+e^{score_{skip}}}$
where, $e_{u},e_{i}\in\mathbb{R}^{n}$ are the user and item embedding, respectively, $\langle\cdot,\cdot\rangle$ is the dot product notation and $score_{skip}$ is a empirically decided hyper-parameter. So, given the score $\text{score}_{item}$ of an item, the user model computes the click likelihood through a softmax function over the recommended item and a predefined skip score. Finally, the user model stochastically selects either click(reward=1) or skip(reward=0) based on the categorical distribution on $p_{item},p_{skip}]$.
Action Representations: Following Jain et al. (2021), we implement continuous item representations sampled from a Gaussian Mixture Model (GMM) with centers around each item category. In this work, we did not use the sub-category in the category system.
11.3 Continuous Control
The MuJoCo (Todorov et al., 2012) benchmarking tasks are a set of standard reinforcement learning environments provided by the MuJoCo physics engine. elow is a brief description of some of the commonly used MuJoCo benchmarking tasks:
Hopper: In the Hopper task, you control a one-legged robot that must learn to hop forward while maintaining balance. The agent needs to find an optimal hopping strategy to maximize forward progress.
Walker2d: This task features a two-legged robot that must learn to walk forward. Similar to the Hopper, the agent must maintain balance while moving efficiently.
HalfCheetah: The HalfCheetah task involves a four-legged cheetah-like robot. The objective is for the robot to learn a coordinated gait that allows it to move forward as rapidly as possible.
Ant: In the Ant task, you control a four-legged ant-like robot. The challenge is for the robot to learn to walk and navigate efficiently through its environment.
11.3.1 Restricted Locomotion in Mujoco
Figure 6 demonstrates "Restricted" locomotion. And here we provide the complete description and justification of the Restricted Mujoco Locomotion tasks below.
Justification: The restricted locomotion setting in Mujoco limits the range of actions the agent is allowed to perform in each dimension. For instance, the wear and tear of an agent’s hardware can easily cause action space to behave like the one visualized in the attached PDF for Hopper. The mixture-of-hypersphere action space is just one way to simulate such asymmetric restrictions. These restrictions apply to the range of torques applied to the joints of hopper and walker, and on the range of forces applied to pendulums.
Complete Description:
@@ 11.3.1 Restricted Locomotion in Mujoco
    Invalid action vectors are replaced with 0. Change to environment’s step function code:
@@
    Invalid action vectors are replaced with -1. Change in code:
@@
    For reference, the action space is $1$-dimensional, with force applied on the cart. The implication is that the cart is pushed in the left direction for $0.02$ (default) seconds. Note that the action vectors are not zeroed because a 0-action is often the optimal action, particularly when the agent starts upright. This would make the learning task trivial, with the optimal strategy being: “learn to select invalid actions”.
Figure 14: Hopper’s 3D visualization of Action Space.
12 Additional Results
12.1 Experiment: Continuous control on Unrestricted Mujoco
Figure 15: TD3 is not suboptimal in Unrestricted Mujoco. We evaluate SAVO against all baselines in the Unrestricted Mujoco continuous control tasks and show that SAVO is competitive with the baselines that already perform optimally. The reason is investigated in §18, where tasks like Inverted Pendulum-v4 and Hopper-v4 have visibly convex Q-landscapes. Thus, SAVO is not expected to significantly outperform TD3 in these benchmarks.
In Mujoco-v4 tasks, the Q-landscape is likely to be easier to optimize than Mujoco-Restricted tasks, and we find that baseline models consistently perform well in all the tasks, unlike Mujoco-Restricted. Based on the performance of SAVO and baselines in Figure 15, we can infer that,
@@ 12.1 Experiment: Continuous control on Unrestricted Mujoco
    The baseline models have sufficient capacity and are well-tuned, as they can solve the standard Mujoco-v4 tasks optimally.
@@
    SAVO performs on par with other methods in Mujoco-v4 tasks where the Q-landscape is easier to optimize.
@@
    Since SAVO outperforms baseline methods only in Mujoco-Restricted, it demonstrates that the reason of SAVO doing better is the presence of a challenging Q-landscape, such as those shown in Figure 1.
12.2 Resetting Baselines
In this section, we clarify the distinction between primacy bias and the challenge of getting stuck in local optima within Q-landscapes. Primacy bias, as addressed in Nikishin et al. (2022); Kim et al. (2024), occurs when an agent is trapped in suboptimal behaviors from early training, and solutions like resetting (reinitializing the parameters of last few layers) and re-learning from the replay buffer mitigate this by avoiding reliance on initially collected samples.
Figure 16: Performance comparisons of Resetting baselines averaged over 5 random seeds, and the seed variance is shown with shading.
However, these methods do not reduce the probability of an actor getting stuck in Q-function local optima (the issue we consider in this work). In fact resetting could cause an otherwise optimal actor to get stuck in suboptima during retraining. To demonstrate this, we conducted a reset baseline experiment, following Nikishin et al. (2022), on TD3 in MineEnv. Here, Full-reset refers to the reset all strategy proposed by Kim et al. (2024), while Last-layer-OOO corresponds to the approach in Nikishin et al. (2022). Finally, TD3 (no reset) represents the standard TD3 algorithm without these extensions. We observed no performance improvements over the standard TD3. In contrast, our method, SAVO, directly addresses this problem by employing an actor architecture specifically designed to navigate non-convex Q-landscapes, making it more robust to local optima.
12.3 Exploration Noise comparison: OUNoise vs Gaussian
Figure 17: OU versus Gaussian Noise. We do not see a significant difference due to this choice, and select OU noise due to better overall performance in experiments
The choice of Ornstein-Uhlenbeck (OU) noise or Gaussian noise for exploration does not make a significant difference and we select OU noise for its better overall performance in initial experiments. This comparison is shown in Figure 17. This finding is consistent with TD3 Fujimoto et al. (2018), which also finds no significant difference between OU and Gaussian noise.
12.4 SAC does not address non-convex Q-landscapes
We compare the performance of SAC, TD3, and TD3 + SAVO across three Mujoco-Restricted tasks. The results (Figure 18) indicate that TD3 + SAVO consistently outperforms the other methods, demonstrating the effectiveness of SAVO in Hopper and Walker2D. In Inverted Pendulum, TD3 + SAVO also shows faster convergence, further highlighting its advantages.
Figure 18: SAC is orthogonal to the effect of SAVO. While SAC is a stochastic extension of TD3 with entropy regularization, SAVO is a plug-in actor architecture that mitigates the challenge of the actor getting stuck in local optima. Thus, tasks where SAC outperforms TD3 differ from tasks where SAVO outperforms TD3. Also, TD3 outperforms SAC in Restricted Hopper and Inverted-Double-Pendulum. However, SAVO+TD3 guarantees improvement over TD3. As we show in §17, SAVO+SAC also mitigates the local optima challenges in SAC.
13 Network Architectures
13.1 Successive Actors
The whole actor has a successive format and each successive actor will receive two pieces of information: the state observation and the action list generated by previous successive actors. Given the concatenation of the input components above, a 4-layer MLP with ReLU will process this information and generate one action for one single successive actor. And this action will be concatenated with the previous action list. After being transformed by an optional action-list-encoder, together with the state information, they become the input of next successive actor’s input. In the end, the action list will be processed with 1-NN to find the nearest discrete action. After this, this action list will be delivered to the selection Q-network.
13.2 Successive Critics
The critic has a one-to-one mapping relationship with the actor. The whole critic consists of a list of successive critics and each successive critic will receive three pieces of information: the state observation, the action list generated by previous successive actors, and the action provided by the corresponding successive actor. Given the concatenation of the input components above, a 2-layer MLP with ReLU will process this information and generate the action’s value for one single successive actor. This value will be used to update itself and the actor with TD-error.
13.3 List Summarizers
In order to extract meaningful information from the list of candidate actions, following Jain et al. (2021) we employed the sequential models and the list-summarizer as follows;
Bi-LSTM: The raw action representations of candidate actions are passed on to the 2-layer MLP followed by ReLU. Then, the output of the MLP is processed by a 2-layer bidirectional LSTM (Huang et al., 2015). Another 2-layer MLP follows this to create the action set summary to be used in the following successive actor.
DeepSet: The raw action representations of candidate actions are passed on to the 2-layer MLP followed by ReLU. Then, the output of the MLP is aggregated by the mean pooling over all the candidate actions to compress the information. Finally, the 2-layer MLP with ReLU provides the resultant action summary to the following successive actor.
Transformer: Similar to the Bi-LSTM variant of the summarizer, we employed the 2-layer MLP with ReLU before inputting the candidate actions into a self-attention and feed-forward network to summarize the information. Afterward the summarization will be part of the input of the following successive actor.
13.4 Feature-wise Linear Modulation (FiLM)
Feature-wise Linear Modulation (Perez et al., 2018), is a technique commonly applied in neural networks for tasks like image recognition. It enhances adaptability by dynamically adjusting intermediate feature representations. Using learned parameters from one layer, FiLM linearly modulates features in another layer, allowing the network to selectively emphasize or de-emphasize aspects of the input data. This flexibility is beneficial for capturing complex and context-specific relationships, improving the model’s performance in various tasks.
13.5 Selection Q-network
The selection Q-network sequentially evaluates the Q-value of the retrieved candidate actions by the cascading actors. Thus, it receives a concatenated information of state and an action embedding for each candidate action. Then, it selects the action with the largest Q-value amongst candidate actions to act on the environment.
14 More experimental results
14.1 More Complex RecSim: Increasing Size of Action Space
We test the robustness of our method to more challenging Q-value landscapes in Figure 19 in Appendix 14.1. In RecSim, we vary the action space size, from $100K$ to $500K$. The results show that SAVO outperforms the baselines, maintaining its robust performance even as the action complexity increases. In contrast, the baselines experienced performance deterioration as action sizes grew larger.
Figure 19: Increasing RecSim action set size: (Left) $100K$ items, (Right) $500K$ items (6 seeds).
14.2 Design Choices: Action summarizers
In the exploration of action summarizer design choices, three key architectures were considered: Deepset, LSTM, and Transformer models, each represented by SAVO, SAVO-lstm, and SAVO-transformer in Fig.20, respectively. In the discrete tasks, the comparison revealed a preference for the deepset architecture over LSTM and Transformer. In the continuous domain, however, the results were rather varied, indicating that the effectiveness of the action summarizer depends on the specific use case. The nuanced differences among these architectures contribute to the complexity of the task, and further research is needed to determine the optimal design for action summarization in both discrete and continuous contexts.
Figure 20: Comparison of action summarizers: the results are averaged over 5 random seeds, and the seed variance is shown with shading.
14.3 Conditioning on previous actions: FiLM v/s MLP
In the examination of conditioning on previous actions, two distinct approaches, Feature-wise Linear Modulation (FiLM) and Multi-Layer Perceptron (MLP), represented by FiLM and non-FiLM variants in Fig.21, were scrutinized for their efficacy. In the discrete tasks, the results unveiled a notable preference for FiLM over non-FiLM implementations, highlighting its effectiveness in leveraging information from prior actions for improved conditioning. However, in the continuous domains, the comparison between FiLM and MLP yielded varied outcomes, suggesting that the choice between these approaches is intricately tied to the specific task context. The nuanced performance differences observed underscore the need for continued research to ascertain the optimal approach for conditioning on previous actions and to enhance model adaptability across diverse applications.
Figure 21: Comparison of how to condition on previous actions: the results are averaged over 5 random seeds, and the seed variance is shown with shading.
Figure 22: Ablation study of SAVO Variations over 5 random seeds shows that every technical component introduced in SAVO contributes to its performance.
14.4 Per-Environment Ablation Results
@@
Removing information about preceding actions does not significantly degrade SAVO’s performance since preceding actions’ Q-values are indirectly incorporated into the surrogates’ training objective (see Eq. 8), except for MineWorld where this information helps improve efficiency.
@@
14.5 Surrogate Approximation Error Analysis
In Figure 23, we analyze the surrogate approximation error across different environments to evaluate how well the surrogate Q-functions approximate the true thresholded Q-function during training. The surrogate error, i.e., the MSE loss from Equation 8, is expressed as a percentage of the Bellman error to measure how closely the surrogate tracks updates to the Q-function. This analysis is important because surrogates aim to simplify optimization while still allowing gradients to propagate effectively.
Low Surrogate Error Across Training. In most environments, the surrogate error converges to a relatively low value between 1–10% of the Bellman error, showing that the surrogates provide a reliable approximation. This indicates that the surrogate functions are well-suited for actor updates, not introducing large errors in the Q-landscape and staying current with new optimal regions. The surrogate error stays consistently low across various tasks, including restricted locomotion (e.g., Hopper, Walker2D) and dexterous manipulation (e.g., Adroit Pen, Adroit Hammer). This demonstrates that the surrogate functions work well across diverse environments with varying levels of complexity.
Non-zero loss shows Smoothness in Flat Regions. The surrogate error remains positive throughout training, including in flat regions of the Q-landscape. This ensures that gradients can still propagate, preventing the actor from getting stuck in areas without gradient information.
Behavior in the Inverted Double Pendulum (Restricted). For the Inverted Double Pendulum (Restricted) environment, the surrogate error increases towards the end of training. This happens because the agent has already converged, and the increase in error reflects overtraining, which is consistent with the observation of an unstable drop in task performance for certain seeds.
Overall, this analysis shows that surrogate functions effectively simplify the Q-value landscape, closely track Q-function updates, and maintain robustness across different tasks, justifying their effectiveness in enabling gradient flow in SAVO. This results in SAVO outperforming the SAVO - Approximation baseline, as shown in Figure 7(b) and Figure 22.
Figure 23: Surrogate Approximation Error Analysis. The plot shows the surrogate approximation error as a percentage of the Bellman error during training across various environments: $\frac{\text{Surrogate Approximation Error}}{\text{Bellman Error}}\%$. In most tasks, the surrogate loss converges to a relatively low value (within 1–10% of the Bellman error), indicating that (i) the surrogates effectively track updates to the Q-function, and (ii) the surrogate loss remains strictly positive, highlighting the smoothness of the surrogate landscape, especially in flat regions, where the exact approximation is undesirable to maintain effective gradient propagation. Notably, for the Inverted Double Pendulum (Restricted) environment, a rise in approximation error is observed towards the end of training. Upon further investigation, this was attributed to overtraining after the agent had already converged, corresponding to an unstable decline in task performance.
14.6 Q-Smoothing Analysis: Discrete vs. Continuous Action Spaces
(a) MineWorld
(b) RecSim
(c) RecSim-Data
(d) Hopper (Restricted)
(e) Walker2D (Restricted)
Figure 24: Impact of Q-smoothing. The plots compare the performance of baselines with and without Q-smoothing. Results are averaged over 5 random seeds, with shading indicating variance. Q-smoothing benefits discrete tasks but has negligible impact in continuous action spaces.
Baseline MineWorld RecSim RecSim-Data Hopper Walker2D
TD3 $0.6$ $4.5$ $42$ $2100$ $2700$
Wolpertinger [Naive] $0.0$ $3.9$ $40$ $\underline{1650}$ $\underline{2900}$
Wolpertinger [Q-Smoothing] $\underline{0.9}$ $\underline{5.0}$ $\underline{46}$ $1850$ $2400$
SAVO (Ours) $\bm{0.98}$ $\bm{5.5}$ $\bm{51}$ $\bm{2500}$ $\bm{3200}$
Table 2: Q-smoothing in discrete tasks. We compare the performance of baselines with and without Q-smoothing across tasks. Underline denotes which variant, naive or Q-smoothing, is used in the paper results. Wolpertinger [Naive] significantly underperforms in discrete action space tasks (denoted in red), and thus, we reported results on Wolpertinger [Q-Smoothing] in the paper. In continuous action space tasks, there was no benefit to Q-smoothing, and thus we chose to report results on Wolpertinger [Naive] as it is closer to the underlying TD3 algorithm. Note that the same Q-smoothing principle is applied for TD3 and SAVO, too, i.e., their Q-function is smoothed for better gradients in discrete action spaces, but unsmoothed Q-function is used in continuous action spaces.
The approximate surrogates introduced in §4.2.2 also have a smoothing effect on the Q-landscape that might ease gradient flow. A similar smoothing can be applied to the primary Q-function. We found such Q-smoothing, which involves learning an auxiliary Q-function to approximate and smooth the primary Q-function, to be essential for discrete action spaces. Q-smoothing facilitates the necessary gradient flow in discrete action space tasks because the primary Q-function is only trained on action representations corresponding to a finite number of discrete actions, while the intermediate action representations might have arbitrary values. By learning an approximate Q-function, the regions between the true action representations are smoothed, facilitating gradient flow.
Thus, in all baselines and SAVO in discrete action space tasks, we included Q-smoothing. However, we did not notice any benefit of Q-smoothing in continuous action space tasks, and thus, all baselines and SAVO do not have Q-smoothing. SAVO still has surrogate smoothing in all environments, because non-smoothed surrogates do not let gradient flow through flat regions.
To demonstrate the impact of Q-smoothing in both discrete and continuous action spaces, we conducted a detailed analysis across several tasks in Figure 24 and Table 2. This section investigates its efficacy and highlights the nuanced differences in its utility across environments.
Discrete Action Spaces: Importance of Q-Smoothing.
For discrete tasks, smoothing the Q-function significantly enhances performance by mitigating the complexity of local optima in diverse Q-value landscapes. This experiment primarily compares 1-Actor k-samples Wolpertinger-Naive and Wolpertinger-Q-smoothing approaches. As shown in Fig. 24, Q-smoothing is essential for Wolpertinger to perform well, while the non-smoothed counterparts significantly suffer in MineWorld and RecSim tasks. Note that the TD3 and SAVO results also include Q-smoothing.
Continuous Action Spaces: Limited Impact of Q-Smoothing.
In continuous action spaces, Q-smoothing does not yield a significant performance gain. In Wolpertinger, both the naive and Q-smoothing variants show comparable performance, indicating sufficient gradient information is present throughout the action space (unlike discrete action space tasks that have missing true Q-values).
For these tasks, as shown in Fig. 24, the introduction of Q-smoothing neither improves nor degrades performance. This justifies its exclusion from our continuous action space experiments and explains why we reported results for Wolpertinger [Naive] in these environments, as it is closer to the underlying TD3 algorithm. Note that the TD3 and SAVO results also exclude Q-smoothing.
Conclusion. Q-smoothing is crucial for discrete action space tasks, as demonstrated by its strong performance in our results. However, it provides no added value for continuous tasks. Consequently, our baselines reflect these observations, ensuring fair comparisons across all evaluated methods.
14.7 Specialized Initialization Strategies for Diversity in SAVO
To explore the potential impact of diverse policy and surrogate value function initializations on algorithm performance, we tested two specialized initialization strategies beyond the default Xavier initialization (Glorot & Bengio, 2010):
  • Xavier (default). Weights are initialized with the default initialization: $w\simXavier$-init
  • Random. Weights are initialized from a standard normal distribution, i.e., $w\sim\mathcal{N}(0,1)$.
  • Add. Weights are initialized using Xavier initialization, followed by the addition of scaled standard normal noise, i.e., $x\sim\text{Xavier-init}$, $y\sim 0.5\cdot\mathcal{N}(0,1)$, and $w=x+y$.
We compare these specialized initialization strategies in various tasks, with reward curves reported in Fig. 25 and summarized below:
  • MineWorld: Add $\approx$ Random $\approx$ Xavier
  • RecSim: Add $\approx$ Random $\approx$ Xavier
  • Hopper (Restricted): Add $\approx$ Random $\approx$ Xavier
  • Adroit Door: Add $\approx$ Random << Xavier
Findings. The results indicate that specialized initialization strategies aimed at increasing diversity do not particularly improve performance. Across most tasks, Add and Random strategies perform similarly to standard Xavier initialization. However, in the Adroit Door task, the specialized initializations underperform compared to Xavier, suggesting that task-specific factors might influence the effectiveness of standard initialization strategies.
Conclusion. While our experiments show no significant benefit from specialized initialization strategies, the idea of explicitly incorporating diversity into the optimization process remains promising. We believe that designing algorithms with explicit diversity objectives throughout training could serve as a valuable heuristic in future work.
(a) MineWorld
(b) RecSim
(c) Hopper (Restricted)
(d) Adroit Door
Figure 25: Specialized Initialization Strategies. Reward curves compare Random and Add strategies to standard Xavier initialization across 4 tasks, showing no significant advantage of specialized initialization for increasing diversity.
15 Experiment Details
15.1 Aggregated Results: Performance Profiles
Agarwal et al. (2021) proposed a robust means to rigorously validate the efficacy of our approach. Through the incorporation of the suggested performance profile, we have conducted a more thorough comparison of our approach against baselines, resulting in a comprehensive understanding of the inherent statistical uncertainty in our results. In Figure 7(a), the x-axis illustrates normalized scores across all tasks, employing min-max scaling to normalize scores based on the initial performance of untrained agents aggregated across random seeds (i.e., Min) and the final performance presented in Figure 8 (i.e., Max).
Figure 7(a) reveals the consistent high performance of our method across various random seeds, with its curve consistently ranking at the top of the x-axis changes, while baseline curves exhibit earlier declines compared to our approach. This visual evidence substantiates the robustness and reliability of our method across different experimental conditions.
15.2 Implementation Details